the sum of the first twenty-one terms of an arithmetic series is 273. The fifth term is 7. What is the 49th term?
Answer by scott8148(6628)

{[a + (a + 20d)] / 2} * 21 = 273 ___ 21a + 210d = 273 ___ a + 10d = 13

a + 4d = 7

subtracting equations ___ 6d = 6 ___ d = 1

substituting ___ a + 4(1) = 7 ___ a = 3

49th term ___ a + 48d = 51